# On Fock covariance for product systems and the reduced Hao–Ng isomorphism problem by discrete actions

### v1

@@ 3.2. Compactly aligned product systems over right LCM semigroups
Right LCM semigroups include as a special case the quasi-lattice ordered semigroups considered in [42]. In particular, a unital right LCM semigroup is a quasi-lattice if and only if $P^{*}=\{e\}$, and in this case a right common least multiple is unique. Artin monoids [6], the Baumslag-Solitar monoids $B(m,n)^{+}$ [21, 38, 49], and the semigroup $R\rtimes R^{\times}$ of affine transformations of an integral domain $R$ that satisfies the GCD condition [37, 43] are examples of right LCM semigroups.
@@
We say that $X$ is compactly aligned if $i_{p}^{w}(k_{p})i_{q}^{w}(k_{q})\in{\mathcal{K}}(X_{w})$ for all $k_{p}\in{\mathcal{K}}(X_{p})$ and $k_{q}\in{\mathcal{K}}(X_{q})$, whenever $pP\cap qP=wP$. It was established in [11] that this definition is independent of the choice of $w$. For a representation $t$ of a product system $X$ we use the notation
@@
Henceforth let us assume that the right LCM semigroup $P$ is in a group $G$.
@@
Let $P$ be a unital right LCM semigroup that is a subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$ in the sense of Fowler. Then $X$ is compactly aligned if and only if for every $p,q\in P$ with $pP\cap qP=wP$ we have that
@@
First suppose that $X$ is compactly aligned, and let $p,q\in P$ with $pP\cap qP=wP$. Let $k_{p}\in{\mathcal{K}}(X_{p})$ and $k_{q}\in{\mathcal{K}}(X_{q})$. By compact alignment we have that $i_{p}^{w}(k_{p})i_{q}^{w}(k_{q})\in{\mathcal{K}}(X_{w})$. Hence for $r\in wP$ we get
@@
On the other hand for $r\not\in wP$, we have that $\lambda^{(w)}(i_{p}^{w}(k_{p})i_{q}^{w}(k_{q}))\xi_{r}=0$. If $r\notin qP$ then $\lambda^{(q)}(k_{q})\xi_{r}=0$. If $r\in qP$ then $\lambda^{(q)}(k_{q})\xi_{r}\in X_{r}$, but then it has to be that $r\notin pP$, and so $\lambda^{(p)}(k_{p})(\lambda^{(q)}(k_{q})\xi_{r})=0$. In all cases, for $r\notin wP$ we have that
@@
there exists a $k_{w}\in{\mathcal{K}}(X_{w})$ such that $\lambda^{(p)}(k_{p})\lambda^{(q)}(k_{q})=\lambda^{(w)}(k_{w})$. For every $\xi_{w}\in X_{w}$ we get that
@@
Therefore we have that $i_{p}^{w}(k_{p})i_{q}^{w}(k_{q})=k_{w}\in{\mathcal{K}}(X_{w})$, and the proof is complete. ∎
@@
In [11, Proposition 2.4] it is shown that, if $w\in P$ and $r\in P^{*}$, then $i_{w}^{wr}(k_{w})\in{\mathcal{K}}(X_{wr})$ and $t^{(wr)}(i_{w}^{wr}(k_{w}))=t^{(w)}(k_{w})\text{ for all }k_{w}\in{\mathcal{K}}(X_{w})$, when $t$ is Nica covariant. Moreover, in the discussion following [11, Definition 2.9] it is shown that
@@
Moreover we have that $\mathbf{K}_{pP,t_{\ast}}=t^{(p)}({\mathcal{K}}(X_{p}))$. Therefore, if =${\mathcal{F}}=\{p_{1}P,\dots,p_{n}P\}$ is a finite $\cap$-closed subset of ${\mathcal{J}}$, then we have that
@@
irrespectively of the choice of the elements $p_{1},\dots,p_{n}$. It is implicit in [11], and proven in a more general context in [23, Theorem 6.4], that the cores $\mathbf{K}_{p_{i}P,\lambda_{\ast}}$ for distinct $p_{i}P$’s are linearly independent in $\mathbf{B}_{{\mathcal{F}},\lambda_{\ast}}$.
It is remarked in [11] that ${\mathcal{N}}{\mathcal{T}}(X)$ admits a coaction since Nica covariance is a graded relation. Let ${\mathcal{N}}X$ be the induced Fell bundle in ${\mathcal{N}}{\mathcal{T}}(X)$. By [11, Proposition 4.3] we have that
@@
for $p_{1},q_{1},\dots,p_{n},q_{n}\in P$ and $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$ such that $q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P=\emptyset$. In particular, we have that $\mathbf{K}_{\emptyset,t_{\ast}}=(0)$.
@@
We first consider the case $\varepsilon=\varepsilon^{\prime}=1$. In this case we have that $q_{n}^{-1}p_{n}\dots q_{1}^{-1}p_{1}P=\emptyset$ and we wish to show that
@@
and so Nica covariance will imply the required identity. In order to reach a contradiction suppose that $p_{1}P\cap q_{1}r_{n-1}P\neq\emptyset$, so there exists a $q\in P$ such that $p:=r_{n-1}^{-1}q_{1}^{-1}p_{1}q\in P$. By the construction of the $r_{k}$ and the $s_{k}$ for $k=1,2,\dots,n-1$ we have that
@@
Therefore for every $k=1,2\dots,n-1$ we have that
@@
Because of the choice of $p=r_{n-1}^{-1}q_{1}^{-1}p_{1}q$ we also have that
@@
Therefore we have that $s_{1}\cdots s_{n-2}s_{n-1}p\in q_{n}^{-1}p_{n}\dots q_{1}^{-1}p_{1}P=\emptyset$ which is a contradiction, and the proof for the first case is complete.
Suppose now that $\varepsilon=1$ and $\varepsilon^{\prime}=0$. We have that $p_{n}q_{n-1}^{-1}\dots q_{1}^{-1}p_{1}P=\emptyset$ and we wish to show that
@@
Note that we must also have that $q_{n-1}^{-1}p_{n-1}\dots q_{1}^{-1}p_{1}P=\emptyset$, and therefore the previous case implies that
@@
Suppose now that $\varepsilon=0$ and $\varepsilon^{\prime}=1$. We have that $q_{n}^{-1}p_{n}\dots p_{2}q_{1}^{-1}P=\emptyset$ and we wish to show that
@@
Finally, suppose that $\varepsilon=\varepsilon^{\prime}=0$. We have that $p_{n}q_{n-1}^{-1}\dots p_{2}q_{1}^{-1}P=\emptyset$ and we wish to show that
@@
We have to show that $t_{\ast}(b_{{\mathbf{x}}})=t^{(w)}\left(k_{w}\right)$. We will use that $t^{(w)}(k_{w})=t_{\ast}(\hat{t}^{(w)}(k_{w}))$. Since ${\mathbf{x}}=qq^{-1}pp^{-1}P$ we have that $b_{{\mathbf{x}}}\in\mathbf{K}_{{\mathbf{x}},\hat{t}_{\ast}}$. Moreover, since ${\mathbf{x}}=wP=ww^{-1}P$ we have that
@@
Fix $r\in\mathbf{x}$. Then we have that
@@
On the other hand, by Proposition 2.5 we have that
@@
there exists a $k_{p}\in{\mathcal{K}}(X_{p})$ such that $q_{\operatorname{c}}^{\rm{N}}(b)=\tilde{t}^{(p)}(k_{p})$. In particular we have that
@@
For condition (ii), first let =${\mathcal{F}}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$ be a $\cap$-closed subset of ${\mathcal{J}}$ where $\mathbf{x}_{i}=p_{i}P$ and $b_{\mathbf{x}_{i}}\in\mathbf{K}_{\mathbf{x}_{i},\hat{t}_{\ast}}$ for $i=1,\dots,n$. Note that without loss of generality we may suppose that ${\mathbf{x}}_{i}\not={\mathbf{x}}_{j}$ for $i\not=j$ and hence the C*-subalgebras $\mathbf{K}_{p_{i}P,\lambda_{\ast}}$’s are linearly independent in $\mathbf{B}_{{\mathcal{F}},\lambda_{\ast}}$, by [23, Theorem 6.4]. Suppose that
@@
From the comments above, for each $i=1,\dots,n$ we may pick a $k_{p_{i}}\in{\mathcal{K}}(X_{p_{i}})$ such that $q_{\operatorname{c}}^{\rm{N}}(b_{\mathbf{x}_{i}})=\tilde{t}^{(p_{i})}(k_{p_{i}})$. Then $\lambda_{\ast}(b_{{\mathbf{x}}_{i}})=\lambda^{(p_{i})}(k_{p_{i}})$ for all $i=1,\dots,n$, and thus for each $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$ we have that
@@
by using Proposition 2.5. Injectivity of $t_{r}$ now yields that $\sum_{i=1}^{n}\lambda^{(p_{i})}(k_{p_{i}})\xi_{r}=0$ for all $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$. On the other hand we have that $\lambda^{(p_{i})}(k_{p_{i}})\xi_{r}=0$ for all $r\not\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$ and $i=1,\dots,n$, and therefore we obtain that
@@
As the $\mathbf{K}_{\bullet}$-cores are linearly independent by [23, Theorem 6.4], we deduce that $\lambda^{(p_{i})}(k_{p_{i}})=0$ for all $i=1,\dots,n$, and thus $k_{p_{i}}=0$ for all $i=1,\dots,n$, from the injectivity of $\lambda^{(p_{i})}$. In particular we have that
@@
Then the previous arguments show that $q_{\operatorname{c}}^{\rm{N}}(b_{{\mathbf{x}}_{i}})=0$ for all $i=1,\dots,n$. Moreover by Proposition 3.5 for $\tilde{t}$ we also have that $q_{\operatorname{c}}^{\rm{N}}(b_{\emptyset})=0$. Hence we get that
@@
As a corollary we have an alternative proof of [11, Proposition 4.3].
@@ 3.3. The product system of a semigroup
One class of product systems of particular interest arises from the semigroup representations of $P$ itself. Fock covariance in this case is described by Laca and Sehnem [34]. We consider $P$ to be a unital subsemigroup of a discrete group $G$. Following the notation of [34], we write $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ for the words of even length where $p_{1},q_{1},\dots,p_{n},q_{n}\in P$. A word $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ is called neutral if $p_{1}^{-1}q_{1}\cdots p_{n}^{-1}q_{n}=e$. For a word $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ we write
@@
Let $w\colon P\to{\mathcal{B}}(H)$ be an isometric representation of $P$ (i.e., every $w_{p}$ is an isometry) and let $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ where $K(\alpha)\not=\emptyset$. Since every $w_{p}$ is an isometry we have that
@@
In particular, if $\alpha$ is neutral we obtain that
@@
In [34, Proposition 3.2] it is shown that if a map $w\colon P\to{\mathcal{B}}(H)$ satisfies
@@
Since conditions (T1)–(T4) are graded with respect to the grading of ${\mathcal{T}}(P)$, the C*-algebra ${\mathcal{T}}_{u}(P)$ is a quotient of ${\mathcal{T}}(P)$ by an induced ideal, and therefore ${\mathcal{T}}_{u}(P)$ admits a coaction by $G$. By using similar arguments as in Proposition 2.7 it can be proved that
@@
and hence combining with Proposition 2.2 we obtain that
@@
In [34, Corollary 3.19, Corollary 3.20] it is shown that there is a *-epimorphism ${\mathcal{T}}_{u}(P)\to\mathrm{C}^{*}_{\lambda}(P)$ that is equivariant and injective on the fixed point algebra. Hence ${\mathcal{T}}_{u}(P)$ coincides with the universal C*-algebra of the Fell bundle induced in $\mathrm{C}^{*}_{\lambda}(P)$; see also [22, Remark 3.15].
@@
by the canonical *-isomorphism induced by $\operatorname{ad}_{W_{P}}$, since $W_{P}\lambda_{p}(v_{p})W_{P}^{*}=V_{p}$ for all $p\in P$. Since this *-isomorphism is canonical we have that the Fell bundles induced in ${\mathcal{T}}_{\lambda}(X_{P})$ and $\mathrm{C}^{*}_{\lambda}(P)$ are isomorphic, and therefore we also have that
@@
induces a bijection between the unital (equivariant) representations of $X_{P}$ and the unital (resp. equivariant) isometric representations of $P$. Hence we have that
@@
by a canonical *-isomorphism. Note that Laca and Sehnem [34, Definition 3.6] coin ${\mathcal{T}}_{u}(P)$ as the Toeplitz algebra of $P$; however we will not use this terminology, as ${\mathcal{T}}_{u}(P)$ is not ${\mathcal{T}}(X_{P})$. Below we give an alternative proof of (3.6) that squares with our characterisation of Fock covariance.
@@
To fix notation, let ${\mathcal{T}}(P)=\mathrm{C}^{*}(v)$, ${\mathcal{T}}(X_{P})=\mathrm{C}^{*}(\hat{t})$ and ${\mathcal{T}}_{u}(P)=\mathrm{C}^{*}(\tilde{u})$. It is readily verified that the association $t\mapsto w_{t}$ is a bijection between the unital equivariant representations of $X_{P}$ and the unital equivariant isometric representations of $P$. When it is clear from the context we will simply write $w$ instead of $w_{t}$. With this notation we have that
@@
For condition (T2), let $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ be a neutral word such that $K(\alpha)=\emptyset$. We then have that
@@
Let $r\in\mathbf{x}$ and $\xi_{r}=\mu v_{r}\in X_{P,r}$ with $\mu\in\mathbb{C}$. Then Proposition 2.5, along with that $\dot{V}_{\alpha}=\dot{V}_{\beta}$, yields
@@
for the unitary $W_{P}$ of (3.4). Hence, we have that
@@
For condition (T4), let $F$ be a finite set of neutral words, and let $\alpha$ be a neutral word such that $K(\alpha)=\bigcup_{\beta\in F}K(\beta)$. Let ${\mathcal{F}}$ be the $\cap$-closure of $\{K(\beta)\mid\beta\in F\}\cup\{K(\alpha)\}\cup\{\emptyset\}$. For each $D\subseteq F$ we write $\beta_{D}$ for the neutral word that arises by concatenating the words $\beta\in D$ in some order. As the induced constructible ideal does not depend on the order of the concatenation we choose (being on neutral words), we have that
@@
Then, for $r\in K(\alpha)$ and $\mu v_{r}\in X_{P,r}$, we have that
@@
where we used that
@@
Note also that in the third equality we used that
@@
In particular, combining (T2) with (T3) yields that $t_{\ast}(\mathbf{K}_{\emptyset,\hat{t}_{\ast}})=(0)$, and thus $t$ satisfies condition (i) of Theorem 3.2.

### v2

@@ 3.2. Compactly aligned product systems over right LCM semigroups
In this section we will only consider group embeddable right LCM semigroups $P$ inside $G$. If the right LCM is unique for every pair of elements $g_{1},g_{2}\in G$, then the pair $(G,P)$ is called quasi-lattice ordered, a setup that was introduced by Nica [46]. Although Nica’s definition is implemented across the group, Exel [19] noted that it can be relaxed to be relevant only for the semigroup, in which case the pair $(G,P)$ is called weakly quasi-lattice ordered. Further examples of right LCM semigroups include the Artin monoids [6], the Baumslag-Solitar monoids $B(m,n)^{+}$ [24, 42, 54], and the semigroup $R\rtimes R^{\times}$ of affine transformations of an integral domain $R$ that satisfies the GCD condition [41, 47].
@@
We say that $X$ is compactly aligned if $i_{p}^{w}(k_{p})i_{q}^{w}(k_{q})\in{\mathcal{K}}(X_{w})$ for all $k_{p}\in{\mathcal{K}}(X_{p})$ and $k_{q}\in{\mathcal{K}}(X_{q})$, whenever $pP\cap qP=wP$. It was established in the discussion following [13, Definition 2.9] that this definition is independent of the choice of $w$. For a representation $t$ of a product system $X$ we use the notation
@@
Let $P$ be a unital right LCM semigroup that is a subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$ in the sense of Fowler. Then $X$ is compactly aligned if and only if for every $p,q\in P$ with $pP\cap qP=wP$ we have
@@
First suppose that $X$ is compactly aligned, and let $p,q\in P$ with $pP\cap qP=wP$. Let $k_{p}\in{\mathcal{K}}(X_{p})$ and $k_{q}\in{\mathcal{K}}(X_{q})$. By compact alignment we have $i_{p}^{w}(k_{p})i_{q}^{w}(k_{q})\in{\mathcal{K}}(X_{w})$. Hence for $r\in wP$ we get
@@
On the other hand for $r\not\in wP$, we have $\lambda^{(w)}(i_{p}^{w}(k_{p})i_{q}^{w}(k_{q}))\xi_{r}=0$. If $r\notin qP$ then $\lambda^{(q)}(k_{q})\xi_{r}=0$. If $r\in qP$ then $\lambda^{(q)}(k_{q})\xi_{r}\in X_{r}$, but then it has to be that $r\notin pP$, and so $\lambda^{(p)}(k_{p})(\lambda^{(q)}(k_{q})\xi_{r})=0$. In all cases, for $r\notin wP$ we have
@@
there exists a $k_{w}\in{\mathcal{K}}(X_{w})$ such that $\lambda^{(p)}(k_{p})\lambda^{(q)}(k_{q})=\lambda^{(w)}(k_{w})$. For every $\xi_{w}\in X_{w}$ we get
@@
Therefore we have $i_{p}^{w}(k_{p})i_{q}^{w}(k_{q})=k_{w}\in{\mathcal{K}}(X_{w})$, and the proof is complete. ∎
@@
In [13, Proposition 2.4] it is shown that, if $w\in P$ and $r\in P^{*}$, then $i_{w}^{wr}(k_{w})\in{\mathcal{K}}(X_{wr})$ and $t^{(wr)}(i_{w}^{wr}(k_{w}))=t^{(w)}(k_{w})\text{ for all }k_{w}\in{\mathcal{K}}(X_{w})$, when $t$ is Nica covariant. Moreover, in the discussion following [13, Definition 2.9] it is shown that
@@
Moreover we have $\mathbf{K}_{pP,t_{\ast}}=t^{(p)}({\mathcal{K}}(X_{p}))$. Therefore, if =${\mathcal{F}}=\{p_{1}P,\dots,p_{n}P\}$ is a finite $\cap$-closed subset of ${\mathcal{J}}$, then we have
@@
irrespectively of the choice of the elements $p_{1},\dots,p_{n}$. It is implicit in [13], and proven in a more general context in [26, Theorem 6.4], that the cores $\mathbf{K}_{p_{i}P,\lambda_{\ast}}$ for distinct $p_{i}P$’s are linearly independent in $\mathbf{B}_{{\mathcal{F}},\lambda_{\ast}}$.
It is shown in [8, Proposition 3.5] that ${\mathcal{N}}{\mathcal{T}}(X)$ admits a coaction since Nica covariance is a graded relation. Let ${\mathcal{N}}X$ be the induced Fell bundle in ${\mathcal{N}}{\mathcal{T}}(X)$. By [13, Proposition 4.3] we have
@@
for $p_{1},q_{1},\dots,p_{n},q_{n}\in P$ and $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$ such that $q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P=\emptyset$. In particular, we have $\mathbf{K}_{\emptyset,t_{\ast}}=(0)$.
@@
We first consider the case $\varepsilon=\varepsilon^{\prime}=1$. In this case we have $q_{n}^{-1}p_{n}\dots q_{1}^{-1}p_{1}P=\emptyset$ and we wish to show that
@@
and so Nica covariance will imply the required identity. In order to reach a contradiction suppose that $p_{1}P\cap q_{1}r_{n-1}P\neq\emptyset$, so there exists a $q\in P$ such that $p:=r_{n-1}^{-1}q_{1}^{-1}p_{1}q\in P$. By the construction of the $r_{k}$ and the $s_{k}$ for $k=1,2,\dots,n-1$ we have
@@
Therefore for every $k=1,2\dots,n-1$ we have
@@
Because of the choice of $p=r_{n-1}^{-1}q_{1}^{-1}p_{1}q$ we also have
@@
Therefore we have $s_{1}\cdots s_{n-2}s_{n-1}p\in q_{n}^{-1}p_{n}\dots q_{1}^{-1}p_{1}P=\emptyset$ which is a contradiction, and the proof for the first case is complete.
Suppose now that $\varepsilon=1$ and $\varepsilon^{\prime}=0$. We have $p_{n}q_{n-1}^{-1}\dots q_{1}^{-1}p_{1}P=\emptyset$ and we wish to show that
@@
Note that we must also have $q_{n-1}^{-1}p_{n-1}\dots q_{1}^{-1}p_{1}P=\emptyset$, and therefore the previous case implies that
@@
Suppose now that $\varepsilon=0$ and $\varepsilon^{\prime}=1$. We have $q_{n}^{-1}p_{n}\dots p_{2}q_{1}^{-1}P=\emptyset$ and we wish to show that
@@
Finally, suppose that $\varepsilon=\varepsilon^{\prime}=0$. We have $p_{n}q_{n-1}^{-1}\dots p_{2}q_{1}^{-1}P=\emptyset$ and we wish to show that
@@
We have to show that $t_{\ast}(b_{{\mathbf{x}}})=t^{(w)}\left(k_{w}\right)$. We will use the equality $t^{(w)}(k_{w})=t_{\ast}(\hat{t}^{(w)}(k_{w}))$. Since ${\mathbf{x}}=qq^{-1}pp^{-1}P$ we have $b_{{\mathbf{x}}}\in\mathbf{K}_{{\mathbf{x}},\hat{t}_{\ast}}$. Moreover, since ${\mathbf{x}}=wP=ww^{-1}P$ we have
@@
Fix $r\in\mathbf{x}$. Then we have
@@
On the other hand, by Proposition 2.5 we have
@@
there exists a $k_{p}\in{\mathcal{K}}(X_{p})$ such that $q_{\operatorname{c}}^{\rm{N}}(b)=\tilde{t}^{(p)}(k_{p})$. In particular we have
@@
For condition (ii), first let =${\mathcal{F}}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$ be a $\cap$-closed subset of ${\mathcal{J}}$ where $\mathbf{x}_{i}=p_{i}P$ and $b_{\mathbf{x}_{i}}\in\mathbf{K}_{\mathbf{x}_{i},\hat{t}_{\ast}}$ for $i=1,\dots,n$. Note that without loss of generality we may suppose that ${\mathbf{x}}_{i}\not={\mathbf{x}}_{j}$ for $i\not=j$ and hence the C*-subalgebras $\mathbf{K}_{p_{i}P,\lambda_{\ast}}$’s are linearly independent in $\mathbf{B}_{{\mathcal{F}},\lambda_{\ast}}$, by [26, Theorem 6.4]. Suppose that
@@
From the comments above, for each $i=1,\dots,n$ we may pick a $k_{p_{i}}\in{\mathcal{K}}(X_{p_{i}})$ such that $q_{\operatorname{c}}^{\rm{N}}(b_{\mathbf{x}_{i}})=\tilde{t}^{(p_{i})}(k_{p_{i}})$. Then $\lambda_{\ast}(b_{{\mathbf{x}}_{i}})=\lambda^{(p_{i})}(k_{p_{i}})$ for all $i=1,\dots,n$, and thus for each $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$ we have
@@
by using Proposition 2.5. Injectivity of $t_{r}$ now yields $\sum_{i=1}^{n}\lambda^{(p_{i})}(k_{p_{i}})\xi_{r}=0$ for all $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$. On the other hand we have $\lambda^{(p_{i})}(k_{p_{i}})\xi_{r}=0$ for all $r\not\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$ and $i=1,\dots,n$, and therefore we obtain
@@
As the $\mathbf{K}_{\bullet}$-cores are linearly independent by [26, Theorem 6.4], we deduce that $\lambda^{(p_{i})}(k_{p_{i}})=0$ for all $i=1,\dots,n$, and thus $k_{p_{i}}=0$ for all $i=1,\dots,n$, from the injectivity of $\lambda^{(p_{i})}$. In particular we have
@@
Then the previous arguments show that $q_{\operatorname{c}}^{\rm{N}}(b_{{\mathbf{x}}_{i}})=0$ for all $i=1,\dots,n$. Moreover by Proposition 3.5 for $\tilde{t}$ we also have $q_{\operatorname{c}}^{\rm{N}}(b_{\emptyset})=0$. Hence we get
@@
As a corollary we have an alternative proof of [13, Proposition 4.3].
@@ 3.3. The product system of a semigroup
One class of product systems of particular interest arises from the semigroup representations of $P$ itself. There is a well-known left regular model in this case considered by Li [41], giving rise to a Fock covariant Fell bundle. The corresponding graded relations were identified by Laca and Sehnem in [38, Section 3], although not explicitly stated like this, see also [25, Remark 3.15]. In this subsection we will show how the Fock covariance description of [38, Section 3] coincides with our general description from Theorem 3.2 in two ways. One is achieved categorically though equation (3.6), and the other is achieved from first principles in Proposition 3.9, thus providing an independent validation of our results.
We consider $P$ to be a unital subsemigroup of a discrete group $G$. Following the notation of [38], we write $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ for the words of even length where $p_{1},q_{1},\dots,p_{n},q_{n}\in P$. A word $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ is called neutral if $p_{1}^{-1}q_{1}\cdots p_{n}^{-1}q_{n}=e$. For a word $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ we write
@@
Let $w\colon P\to{\mathcal{B}}(H)$ be an isometric representation of $P$ (i.e., every $w_{p}$ is an isometry) and let $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ where $K(\alpha)\not=\emptyset$. Since every $w_{p}$ is an isometry we have
@@
In particular, if $\alpha$ is neutral we obtain
@@
In [38, Proposition 3.2] it is shown that, if a map $w\colon P\to{\mathcal{B}}(H)$ satisfies
@@
Since conditions (T1)–(T4) are graded with respect to the grading of ${\mathcal{T}}(P)$, the C*-algebra ${\mathcal{T}}_{u}(P)$ is a quotient of ${\mathcal{T}}(P)$ by an induced ideal, and therefore ${\mathcal{T}}_{u}(P)$ admits a coaction by $G$. By using similar arguments as in Proposition 2.7 it can be proved that we have
@@
and hence combining with Proposition 2.2 we obtain
@@
In [38, Corollary 3.19, Corollary 3.20] it is shown that there is a *-epimorphism ${\mathcal{T}}_{u}(P)\to\mathrm{C}^{*}_{\lambda}(P)$ that is equivariant and injective on the fixed point algebra. Hence ${\mathcal{T}}_{u}(P)$ coincides with the universal C*-algebra of the Fell bundle induced in $\mathrm{C}^{*}_{\lambda}(P)$; see also [25, Remark 3.15].
@@
by the canonical *-isomorphism induced by $\operatorname{ad}_{W_{P}}$, since $W_{P}\lambda_{p}(v_{p})W_{P}^{*}=V_{p}$ for all $p\in P$. Since this *-isomorphism is canonical we have that the Fell bundles induced in ${\mathcal{T}}_{\lambda}(X_{P})$ and $\mathrm{C}^{*}_{\lambda}(P)$ are isomorphic, and therefore we also have
@@
induces a bijection between the unital (equivariant) representations of $X_{P}$ and the unital (resp. equivariant) isometric representations of $P$. Hence we have
@@
by a canonical *-isomorphism. Note that Laca and Sehnem [38, Definition 3.6] coin ${\mathcal{T}}_{u}(P)$ as the Toeplitz algebra of $P$; however we will not use this terminology, as ${\mathcal{T}}_{u}(P)$ is not ${\mathcal{T}}(X_{P})$. Below we give an alternative proof of (3.6) that squares with our characterisation of Fock covariance.
@@
To fix notation, let ${\mathcal{T}}(P)=\mathrm{C}^{*}(v)$, ${\mathcal{T}}(X_{P})=\mathrm{C}^{*}(\hat{t})$ and ${\mathcal{T}}_{u}(P)=\mathrm{C}^{*}(\tilde{u})$. It is readily verified that the association $t\mapsto w_{t}$ is a bijection between the unital equivariant representations of $X_{P}$ and the unital equivariant isometric representations of $P$. When it is clear from the context we will simply write $w$ instead of $w_{t}$. With this notation we have
@@
For condition (T2), let $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ be a neutral word such that $K(\alpha)=\emptyset$. We then have
@@
Let $r\in\mathbf{x}$ and $\xi_{r}=\mu v_{r}\in X_{P,r}$ with $\mu\in\mathbb{C}$. Then Proposition 2.5, along with the equality $\dot{V}_{\alpha}=\dot{V}_{\beta}$, yields
@@
for the unitary $W_{P}$ of (3.4). Hence, we have
@@
For condition (T4), let $F$ be a finite set of neutral words, and let $\alpha$ be a neutral word such that $K(\alpha)=\bigcup_{\beta\in F}K(\beta)$. Let ${\mathcal{F}}$ be the $\cap$-closure of $\{K(\beta)\mid\beta\in F\}\cup\{K(\alpha)\}\cup\{\emptyset\}$. For each $D\subseteq F$ we write $\beta_{D}$ for the neutral word that arises by concatenating the words $\beta\in D$ in some order. As the induced constructible ideal does not depend on the order of the concatenation we choose (being on neutral words), we have
@@
Then, for $r\in K(\alpha)$ and $\mu v_{r}\in X_{P,r}$, we have
@@
where we used the equalities
@@
Note also that in the third equality we used that we have
@@
In particular, combining (T2) with (T3) yields $t_{\ast}(\mathbf{K}_{\emptyset,\hat{t}_{\ast}})=(0)$, and thus $t$ satisfies condition (i) of Theorem 3.2.